# PD-TPE: Parallel Decoder with Text-guided Position Encoding for 3D Visual Grounding

### v1

@@ Abstract
3D visual grounding aims to locate the target object mentioned by free-formed natural language descriptions in 3D point cloud scenes. Most previous work requires the encoder-decoder to simultaneously align the attribute information of the target object and its relational information with the surrounding environment across modalities. This causes the queries’ attention to be dispersed, potentially leading to an excessive focus on points irrelevant to the input language descriptions. To alleviate these issues, we propose PD-TPE, a visual-language model with a double-branch decoder. The two branches perform proposal feature decoding and surrounding layout awareness in parallel. Since their attention maps are not influenced by each other, the queries focus on tokens relevant to each branch’s specific objective. In particular, we design a novel Text-guided Position Encoding method, which differs between the two branches. In the main branch, the priori relies on the relative positions between tokens and predicted 3D boxes, which direct the model to pay more attention to tokens near the object; in the surrounding branch, it is guided by the similarity between visual and text features, so that the queries attend to tokens that can provide effective layout information. Extensive experiments demonstrate that we surpass the state-of-the-art on two widely adopted 3D visual grounding datasets, ScanRefer and NR3D, by 1.8% and 2.2%, respectively. Codes will be made publicly available.
@@ Introduction
Multi-modal learning is the key to enhancing artificial intelligence’s comprehension and awareness of the real world, where the most fundamental and important modalities are visual and language. Visual Grounding (VG), an emerging vision-language task, requires a model to identify and locate objects in scenes described by natural languages. Although Visual Grounding for 2D images has achieved remarkable results, the sparse and complex structure of 3D point clouds, combined with the linguistic diversity introduced by 3D spatial contexts, presents significant challenges for 3D Visual Grounding.
@@
With significant advances in single-modal 3D object detection tasks, the main objective of the 3D Visual Grounding task is to address the issue of ”correct classification but incorrect localization”, which means enabling the model to better understand the spatial layout of the 3D scene. Some models (Wang et al. 2023; Shi, Wu, and Lee 2024) attempt to process visual tokens to better perceive the spatial relationships between detected objects in the scene. Some others (Luo et al. 2022; Wu et al. 2023; Guo et al. 2023) focus on processing the natural language descriptions, using techniques such as text decoupling and referential order awareness to achieve positional and semantic cross-modal alignment.
@@
Figure 1: Comparison between the serial network architecture (a) and our parallel network architecture (b). For grounding results shown in scenes, yellow represents the ground-truth boxes, red is the wrongly localized box, and green is the successfully localized one. Different types of PE act as guidance for the attention maps.
@@
However, existing methods are primarily facing two unresolved issues. On the one hand, the features of the objects themselves, such as shapes and colors, are coupled with the captured spatial relationship features. The former requires encoder-decoders to pay more attention to points near the predicted boxes, while the latter asks the attention map to focus more on tokens forming effective spatial relationships with the target object. For instance, as shown in Fig. 1 (a), the coupled features lead to a dispersed distribution of attention in the attention map, causing the model to locate objects of the same category but in different positions. On the other hand, information from visual and language modalities fails to mutually understand each other. When generating attention maps, it’s challenging for existing methods to directly identify which visual tokens are relevant to the target’s surrounding layout described in the text, leading to the localization of wrong objects.
To address these issues, we introduce an advanced framework named PD-TPE. A double-branch decoder is employed to perform proposal feature decoding and surrounding layout awareness in parallel, which allows the decoder to focus equally on the features of the query proposals themselves and their relationships with other objects in the 3D scene. For text features, the semantically decoupled components of sentences divide the text tokens into main target descriptions and surrounding layout descriptions, serving the main branch and the surrounding branch separately. For visual features, we design a novel Text-guided Relative Position Encoding method, which varies between the two branches. In the main branch, we compute position encoding for each visual token based on its relative position to the 3D boxes predicted by the queries, thus providing clear information to guide the model to focus on the target object itself. In the surrounding branch, we use the similarity between visual and text features to dominate the encoding process, guiding the queries to attend to tokens related to the spatial layout mentioned by the textual information, which is shown in Fig. 1 (b).
@@
Our main contributions can be summarized as follows: (a) We design a parallel decoder with two branches, each holding different attention maps to focus separately on target objects and surrounding layouts; (b) We propose the Text-guided Position Encoding method for the two branches of the decoder, which fully utilizes multimodal inputs to guide the module in focusing on tokens that carry effective information; (c) Extensive experiments show that our PD-TPE delivers new state-of-the-art performance on both ScanRefer and Nr3D benchmarks.
@@
Visual and Language modalities, the two most crucial and fundamental aspects for the machines to comprehend and interact with the 3D real world, give rise to various 3D Vision-Language tasks. 3D Dense Captioning (3DDC) (Chen, Chang, and Nießner 2020; Achlioptas et al. 2020; Jain et al. 2022; Wu et al. 2023) requires a model to accurately localize all objects in a complex 3D scene and generate descriptive captions for them. transformer 3D Visual Grounding (3DVG) (Chen et al. 2020; Yuan et al. 2022; Chen et al. 2023) takes 3D point clouds and language descriptions as inputs and generates corresponding bounding boxes to localize target objects. 3D Question Answering (3DQA) (Azuma et al. 2022; Ma et al. 2023) provides visual information from the 3D scenes for models, and the goal is to answer given textual questions about the 3D scene. All the above tasks mainly focus on the alignment of visual and language features, including the attribute features of objects such as color and shape, as well as the relationships between target objects and their surrounding environment. In this work, we focus on the 3D Visual Grounding tasks (3DVG), attempting to enable the machine to understand inputs in both 3D point clouds and natural language modalities simultaneously.
@@
Previous works can be divided into two-stage methods and one-stage methods according to the overall model structure. The two-stage methods (Chen, Chang, and Nießner 2020; Zhao et al. 2021; Yang et al. 2021; Huang et al. 2022; Guo et al. 2023; Luo et al. 2022) first separately parse the language inputs and perform 3D point cloud object detection. The features of the extracted visual proposals and language queries are fused in the second stage, which facilitates the location of the best-matching objects. Graph-based methods are widely used in early works (Achlioptas et al. 2020; Huang et al. 2021) to represent and process the spatial relationships between the detected proposals, achieving localization in 3D scenes. Recent works (Zhao et al. 2021; Wang et al. 2023; Yang et al. 2023; Zhang, Gong, and Chang 2023; Shi, Wu, and Lee 2024) employ transformers(Vaswani et al. 2017) as the crucial module, finishing the modality alignment and feature abstraction with the cross-attention mechanism.
The one-stage methods are proposed (Jain et al. 2022; Wu et al. 2023; Luo et al. 2022) because, in the two-stage method, objects that are mistakenly overlooked during the pre-training stage cannot be learned or corrected during the fusion stage. (Jain et al. 2022) grounds referential utterances in point clouds by both top-down language guidance and bottom-up objectness guidance. (Luo et al. 2022) incrementally selects key points under language guidance and directly locates target objects with a goal-oriented mining module. (Wu et al. 2023) explicitly decouples textual attributes in sentences and achieves dense alignment between language and point clouds through positional and semantic features. These methods have achieved impressive results, but their limitation is that learning the features of the target objects and their surrounding layout at the same time would make the calculation of the attention map confusing. To this end, we design a one-stage model with a parallel decoder, employing the position encoding method to guide the generation of the attention map and finish precise grounding.
Figure 2: Architecture of our 3D Visual Grounding Framework. With the inputs of natural languages and 3D point clouds, our model consists of a text decoupling module, three encoders for different modalities, a double-branch cross encoder, and a novel parallel decoder. Position Encoding methods are added to cross-attention modules for both main and surrounding branches. The final visual and text output features are aligned with each other to generate the detected boxes.
@@
Problem Formulation
The objective of 3D visual grounding is to accurately localize objects in 3D point cloud scenes according to the given natural language descriptions. In this task, our model takes an input point cloud denoted as $P=\left[p_{in};f_{in}\right]\in\mathbb{R}^{N_{in}\times\left(3+6\right)}$, which represents an indoor 3D scene. Here, $N_{in}$ is the number of points, $p_{in}\in\mathbb{R}^{N_{in}\times 3}$ is the 3D coordinates of each point, and $f_{in}\in\mathbb{R}^{N_{in}\times 6}$ is additional feature for each point, 3 dimensions for localization (XYZ) and 3 dimensions for color (RGB). The input of texts is denoted as $T_{in}$ with length $L$. The output is a set of boxes $B_{out}\in\mathbb{R}^{K_{out}\times 6}$, which represents an estimation of $K_{out}$ corresponding objects with 3 dimensions of box centers (XYZ) and 3 dimensions of box sizes (length-width-height).
Framework Setup
The architecture of our framework is shown in Fig.2. Each module is illustrated in detail in this section.
Input Modal Tokenization.
The input text and 3D point clouds are encoded by the pre-trained RoBERTa (Liu et al. 2019) and PointNet++ (Qi et al. 2017) and produce text tokens $T_{0}\in\mathbb{R}^{L\times D}$ and visual tokens $V_{0}\in\mathbb{R}^{N\times D}$, where $D$ is the dimension of visual and text features and $N$ is the number of visual tokens. Additionally, the GroupFree (Liu et al. 2021) detector is optionally used to detect 3D boxes according to (Wu et al. 2023), which are subsequently encoded as box tokens $B\in\mathbb{R}^{K\times D}$. At the same time, text inputs are decoupled into five semantic components following (Wu et al. 2023), using the off-the-shelf tool from (Schuster et al. 2015; Wu et al. 2019).
@@
We employ BUTD-DETR’s (Jain et al. 2022) cross-encoder module for the intermodulation of visual and text features. Two branches of cross attention separately take one modulation of feature as query and another modulation of feature as key and value. The outputs of this module are represented as text features $T\in\mathbb{R}^{L\times D}$ and visual features $V\in\mathbb{R}^{N\times D}$. Then we select top-K object candidate points from $V$ through their predicted confidence scores, which are defined as $V_{q}\in\mathbb{R}^{K\times D}$ and $K$ represents the number of selected query proposal features.
Parallel Decoder with Main Branch and Surrounding Branch.
We design a decoder with two parallel branches, one is used to extract the main features of query proposals, and another is used to perceive the surrounding environmental layout around each of them. Specifically, the query proposal features $V_{q}$ firstly pass a self-attention layer and act as the query of each cross-attention layer in each branch.
For the text features $T$, we divide the text into main features $T_{m}$ and surrounding features $T_{s}$ based on the decoupled five semantic components. For instance, in “there is a dark brown wooden and leather chair. placed in the table of the kitchen .”, the Main object - “chair” and the Attributes - “dark brown wooden and leather” are main features $T_{m}$ and other parts Auxiliary object, Pronoun, and Relationship are surrounding features $T_{s}$. They would separately act as the key and value of the text cross-attention layer in two branches.
The visual features $V$ act as the key and value of the visual cross-attention layer in both two branches. Inspired by (Shen et al. 2024), we design different novel position encoding methods for each branch, which would be described in detail in Sec. Position Encoding Position Encoding.
Finally, visual-guided and text-guided features in each branch are fused together first, and then features generated from the two branches are fused. The query proposal features are updated as $V^{h}_{q}\in\mathbb{R}^{K\times D}$, where $h$ represents the $h-th$ layer of the decoder.
@@
Following (Wu et al. 2023), the query proposal features $V^{h}_{q}$ in each layer are projected into position $V^{h}_{l}\in\mathbb{R}^{K\times D_{l}}$ and object semantic features $V^{h}_{o}\in\mathbb{R}^{K\times D_{o}}$ for losses calculation, where $D_{l}$ and $D_{o}$ represents the dimensions of two types of features. A box prediction head (Liu et al. 2021) is employed to generate bounding boxes $B_{out}\in\mathbb{R}^{K_{out}\times 6}$.
Figure 3: Different methods to calculate the relevant position between token A and the predicted box: (a) Vertex Relative Positioning (Shen et al. 2024); (b) Center Relative Positioning; (c)(d) Box-surface Relative Positioning for token outside/inside the box.
Position Encoding
Relative Position Encoding for Attention Map.
Position encoding (PE) can effectively refine the attention map, which is crucial for understanding the contextual information of tokens in 3D scenes. For the 3D visual grounding task, we update the calculation method of the attention map for the decoder as follows:
@@
We encode the relative positions of visual tokens and coarsely detected target objects, using grid-sample with a pre-defined table $\mathcal{T}$ to reduce memory overhead:
${\mathcal{E}}=\operatorname{grid\_sample}(\operatorname{MLP}(\mathcal{T}),\;\operatorname{F}(\Delta\mathcal{E})),$ (2)
where $\operatorname{F}(\cdot)$ is a non-linear function and $\Delta\mathcal{E}\in\mathbb{R}^{K\times N\times 3}$ is the calculated relative position. For a detailed explanation of this part, please refer to (Shen et al. 2024).
@@
To represent the relative position between $N$ tokens and $K$ boxes, we find the point on the box surface closest to each token and use the offsets in three dimensions as $\Delta\mathcal{E}=\{(e^{ij}_{x},e^{ij}_{y},e^{ij}_{z})|\quad i=0,1,...,K;j=0,1,...,N\}$, which is shown in Fig.3.
For a box $b=(x_{c},y_{c},z_{c},l,w,h)$ represented by center $C$ and size and a token with position coordinate outside the box $A=(x_{0},y_{0},z_{0})$ as Fig.3(c), the relative coordinate of $A$ can be represented as $\widehat{A}=(\widehat{x},\widehat{y},\widehat{z})=(x_{0}-x_{c},y_{0}-y_{c},z_{0}-z_{c})$. The relative coordinate of a point on the box surface can be represented as $M=(x,y,z)$, so the distance between $A$ and $M$ is:
$MA^{2}=(\widehat{x}-x)^{2}+(\widehat{y}-y)^{2}+(\widehat{z}-z)^{2},$ (3)
where $|x|\geq\frac{l}{2},|y|\geq\frac{w}{2},|z|\geq\frac{h}{2}$. Due to the symmetry of relative positions, we only need to consider the positive half-axis directions in three dimensions, which corresponds to $\frac{1}{8}$ of the entire relative coordinate system. The problem becomes finding the minimum value of the following equation:
$MA^{2}=(|\widehat{x}|-x)^{2}+(|\widehat{y}|-y)^{2}+(|\widehat{z}|-z)^{2}.$ (4)
The extreme value is obtained at the critical point, which means that
$\begin{split}MA^{2}_{min}=(max(|\widehat{x}|-\frac{l}{2},0))^{2}+(max(|\widehat{y}|-\frac{h}{2},0))^{2}\\ +(max(|\widehat{y}|-\frac{w}{2},0))^{2}.\end{split}$ (5)
Therefore, the three-dimension offset can be represented as:
$\begin{split}(e^{ij}_{x},e^{ij}_{y},e^{ij}_{z})=(max(|\widehat{x}|-\frac{l}{2},0),\\ max(|\widehat{y}|-\frac{h}{2},0),max(|\widehat{y}|-\frac{w}{2},0)),\end{split}$ (6)
where $i=0,1,...,K$ and $j=0,1,...,N$.
When $(|\widehat{x}|<\frac{l}{2})\wedge(|\widehat{y}|<\frac{w}{2})\wedge(|\widehat{z}|<\frac{h}{2})$, the token $A^{\prime}=(x^{\prime}_{0},y^{\prime}_{0},z^{\prime}_{0})$ is inside the box as Fig.3(d). The relative coordinate can be represented as $\widehat{A}^{\prime}=(\widehat{x}^{\prime},\widehat{y}^{\prime},\widehat{z}^{\prime})=(x^{\prime}_{0}-x_{c},y^{\prime}_{0}-y_{c},z^{\prime}_{0}-z_{c})$. Only considering the $\frac{1}{8}$ positive area, the minimum relative distance can be written as:
$MA^{\prime}_{min}=\operatorname{min}(\frac{l}{2}-|\widehat{x}^{\prime}|,\frac{w}{2}-|\widehat{y}^{\prime}|,\frac{h}{2}-|\widehat{z}^{\prime}|).$ (7)
The three-dimension offset for tokens inside boxes can be represented as:
$e^{ij}_{o}=\left\{\begin{array}[]{ll}-MA^{\prime}_{min}\ ,&if\frac{d}{2}-|\widehat{o}^{\prime}|=MA^{\prime}_{min}\\ 0\ ,&otherwise\end{array}\right.$ (8)
where $(o,d)=(x,l),(y,w),and(z,h)$. The offset takes the negative of the distance because positive and negative distances represent different relative positions of the tokens.
@@
Compared to directly using the offset between the token and the box center as the relative position shown in Fig.3(b), our Box-surface Relative Positioning method incorporates both relative distance information and the shape information of the box. In contrast to the Vertex Relative Positioning method proposed by (Shen et al. 2024) shown in Fig.3(a), our method requires encoding only once in each layer of cross-attention, making it more time-efficient and memory-saving. For detailed ablation experiments, please refer to Sec. Experiments.
Figure 4: The illustration of Text-guided Position Encoding method. The similarity between visual tokens and surrounding text tokens revises the attention map.

### v2

# PD-APE: A Parallel Decoding Framework with Adaptive Position Encoding for 3D Visual Grounding

@@ Abstract
3D visual grounding aims to identify objects in 3D point cloud scenes that match specific natural language descriptions. This requires the model to not only focus on the target object itself but also to consider the surrounding environment to determine whether the descriptions are met. Most previous works attempt to accomplish both tasks within the same module, which can easily lead to a distraction of attention. To this end, we propose PD-APE, a dual-branch decoding framework that separately decodes target object attributes and surrounding layouts. Specifically, in the target object branch, the decoder processes text tokens that describe features of the target object (e.g., category and color), guiding the queries to pay attention to the target object itself. In the surrounding branch, the queries align with other text tokens that carry surrounding environment information, making the attention maps accurately capture the layout described in the text. Benefiting from the proposed dual-branch design, the queries are allowed to focus on points relevant to each branch’s specific objective. Moreover, we design an adaptive position encoding method for each branch respectively. In the target object branch, the position encoding relies on the relative positions between seed points and predicted 3D boxes. In the surrounding branch, the attention map is additionally guided by the confidence between visual and text features, enabling the queries to focus on points that have valuable layout information. Extensive experiments demonstrate that we surpass the state-of-the-art on two widely adopted 3D visual grounding datasets, ScanRefer and Nr3D.
@@ Introduction
Multi-modal learning is the key to enhancing artificial intelligence’s comprehension and awareness of the real world, where the most fundamental and important modalities are visual and language. Visual Grounding (VG), an emerging vision-language task, requires a model to identify and locate target objects in scenes based on natural language descriptions. Although Visual Grounding for 2D images has achieved remarkable results, the sparse and complex structure of 3D point clouds, combined with the linguistic diversity introduced by 3D spatial contexts, presents significant challenges for 3D Visual Grounding.
@@
With significant advances in single-modal 3D object detection, the main objective of the 3D Visual Grounding task is to address the issue of “correct classification but incorrect localization”, which requires the model to better understand the spatial layout of the 3D scene. Some methods (Wang et al. 2023; Shi, Wu, and Lee 2024) attempt to process visual features to better perceive the spatial layouts among detected objects in the scene. Others (Luo et al. 2022; Wu et al. 2023; Guo et al. 2023) focus on processing the natural language descriptions, using techniques such as text decoupling and referential order awareness to achieve positional and semantic cross-modal alignment.
@@
Figure 1: Comparison between the serial decoder structure (a) and our parallel decoder structure (b). For grounding results shown in scenes, yellow represents the ground-truth boxes, red is the wrongly localized box, and green is the successfully localized one.
@@
However, existing methods have two problems. First, object attribute (e.g., shapes and colors) features and spatial layout features are entangled. The former requires the network to pay more attention to points near the predicted boxes, while the latter asks the attention map to focus more on points forming spatial relationships with target objects. Previous serial structures cannot disentangle different types of attention. For instance, as shown in Fig. 1 (a), the coupled features lead to a dispersed distribution of attention in the attention map, causing the model to locate objects of the same category but in different positions. Second, the spatial information from text tokens is not employed in the visual cross-attention block. Without the guidance of text information, queries can only roughly learn the features of neighboring points. This causes queries to focus on redundant and irrelevant layout information, leading to a dispersion of attention.
To address these problems, we introduce a new framework named PD-APE. A dual-branch decoder is employed to perform target object feature decoding and surrounding layout awareness in parallel. Specifically, the target object branch refines the attention map through local visual features and text features that describe the target’s attributes (e.g. category and color); the surrounding branch captures the relationship between the target and surrounding objects through contextual text tokens and corresponding visual features. Taking the text information into account, we semantically divide the tokens into target descriptions and layout descriptions, serving the target object branch and the surrounding branch, respectively. Moreover, we design an adaptive position encoding method to enhance associated visual features. In the target object branch, we obtain position encoding for each seed point by using its relative position to the surfaces of the predicted boxes. It provides explicit information to guide the model to focus on the target object itself. In the surrounding branch, we formulate the confidence between visual and text features as a gate to dominate the encoding process. This gate guides the queries to focus on points that are related to the spatial layout described by the textual inputs. The attention maps generated by our new framework are shown in Fig. 1 (b).
@@
Our main contributions can be summarized as follows: (a) We design a parallel decoder with two branches, each holding different attention maps to focus separately on target objects and surrounding layouts; (b) We propose an adaptive position encoding method for the two branches of the decoder, which fully utilizes multi-modal inputs to guide the module to focus on points that carry effective information; (c) Extensive experiments show that our PD-APE delivers new state-of-the-art performance on both ScanRefer and Nr3D benchmarks.
@@
Vision and language, the two most essential and foundational modalities for machines to understand and interact with the 3D real world, give rise to various 3D vision-language tasks. 3D Dense Captioning (3DDC) (Chen et al. 2020; Yuan et al. 2022; Chen et al. 2023) requires a model to accurately localize all objects in a complex 3D scene and generate descriptive captions for them. 3D Visual Grounding (3DVG) (Chen, Chang, and Nießner 2020; Achlioptas et al. 2020; Jain et al. 2022) takes 3D point clouds and language descriptions as inputs and generates corresponding bounding boxes to localize target objects. 3D Question Answering (3DQA) (Azuma et al. 2022; Ma et al. 2023) provides visual information from the 3D scenes for models, aiming to answer textual questions about the 3D scene. All of the above tasks primarily focus on aligning visual and language features, encompassing object attributes (e.g., shapes and colors) and the relationships between targets and their environment. In this work, we focus on 3D Visual Grounding tasks (3DVG), enabling machines to simultaneously comprehend inputs from 3D point clouds and natural language.
@@
Previous works can be categorized into two-stage and one-stage methods based on their overall model structure.
The two-stage methods (Chen, Chang, and Nießner 2020; Zhao et al. 2021; Yang et al. 2021; Huang et al. 2022; Guo et al. 2023; Luo et al. 2022) first separately parse the language inputs and perform 3D point cloud object detection. In the second stage, the features of the extracted visual proposals and language tokens are fused, facilitating the identification of the best-matching objects. Graph-based methods are widely used in early works (Achlioptas et al. 2020; Huang et al. 2021) to represent and process the spatial relationships between the detected proposals, achieving localization in 3D scenes. Recent works (Zhao et al. 2021; Wang et al. 2023; Yang et al. 2023; Zhang, Gong, and Chang 2023; Shi, Wu, and Lee 2024) employ transformers (Vaswani et al. 2017) as the key module, leveraging the cross-attention mechanism to achieve modality alignment and feature abstraction.
The one-stage methods (Jain et al. 2022; Wu et al. 2023; Luo et al. 2022) are proposed because, in the two-stage method, objects mistakenly overlooked during the pre-training stage cannot be learned or corrected during the fusion stage. (Jain et al. 2022) grounds referential utterances in point clouds by both top-down language guidance and bottom-up object guidance. (Luo et al. 2022) incrementally selects key points under language guidance and directly locates target objects with a goal-oriented mining module. (Wu et al. 2023) explicitly decouples textual attributes in sentences and achieves dense alignment between language and point clouds through positional and semantic features. These methods have achieved impressive results, but simultaneously learning the features of target objects and their surrounding layout can lead to a dispersed distribution of attention in the attention map. To this end, we design a one-stage model with a parallel decoder, employing the position encoding method to guide the generation of the attention map and achieve precise grounding.
Figure 2: Architecture of our 3D Visual Grounding Framework. With the inputs of natural languages and 3D point clouds, our model consists of three encoders for different modalities, a cross encoder, and a novel parallel decoder. Adaptive Position Encoding methods are added to cross-attention modules for both the target object branch and the surrounding branch. The final visual and text output features are aligned with each other to generate the detected boxes.
@@
Overview
The objective of 3D visual grounding is to accurately localize objects in 3D point cloud scenes according to the given natural language descriptions. In this task, our model uses point cloud sampling to $N_{in}$ points as input, which includes 3 dimensions for localization (XYZ) and 3 dimensions for color (RGB). The input of texts is denoted as $T_{in}$ with length $L$. The output is a set of boxes $B_{out}\in\mathbb{R}^{K_{out}\times 6}$, which represents an estimation of $K_{out}$ corresponding objects with 3 dimensions of box centers (XYZ) and 3 dimensions of box sizes (length-width-height).
The architecture of our framework is shown in Fig. 2. Visual and text inputs first go through their respective modality encoders, then together pass through a cross encoder and a parallel decoder for target object features and surrounding features, and finally complete grounding through dense alignment following (Wu et al. 2023). Each module is illustrated in detail in this section.
Modality Encoders.
The input texts are encoded by the pre-trained RoBERTa (Liu et al. 2019), which produces text tokens with features $T_{0}\in\mathbb{R}^{L\times D}$ and $D$ is the dimension of features. For point cloud inputs, we provide two different backbone architectures: (a) PointNet++ (Qi et al. 2017); (b) a sparse 3D modification of ResNet34 (He et al. 2015) following FCAF3D (Rukhovich, Vorontsova, and Konushin 2022). The former is adopted for a fair comparison with the recent methods, while the latter delivers better performance. The visual encoder samples $N$ seed points with features $V_{0}\in\mathbb{R}^{N\times D}$ from the point clouds. Additionally, the GroupFree (Liu et al. 2021) detector is optionally used to detect 3D boxes according to (Wu et al. 2023), which are subsequently encoded as box tokens $B\in\mathbb{R}^{K\times D}$.
@@
We employ BUTD-DETR’s (Jain et al. 2022) cross-encoder module for the intermodulation of visual and text features. Two branches with cross-attention blocks separately take features of one modality as query and features of another modality as key and value. The outputs of this module are represented as text token features $T\in\mathbb{R}^{L\times D}$ and visual features for seed points (seed point features) $V\in\mathbb{R}^{N\times D}$. Then we select top-K candidates from seed points as queries through their predicted confidence scores, whose features are defined as $V_{q}\in\mathbb{R}^{K\times D}$.
Parallel Decoder with Target Object Branch and Surrounding Branch.
We design a decoder with two parallel branches. One is used to extract the target object features of query proposals, and another is used to perceive the environmental layout around each of them. The query proposal features $V_{q}$ firstly pass a self-attention layer and act as the query of each cross-attention layer in each branch.
For the text features $T$, we divide the text tokens into five semantic components following (Wu et al. 2023), using the off-the-shelf tool from (Schuster et al. 2015; Wu et al. 2019). Based on the five above components, target object tokens with features $T_{m}$ serve for the target object branch, and other tokens with features $T_{s}$ serve for the surrounding branch. For instance, in “there is a dark brown wooden chair. placed in the table of the kitchen .”, the Main object - “chair” and the Attributes - “dark brown wooden” are target object features $T_{m}$ and other parts Auxiliary object, Pronoun, and Relationship are surrounding features $T_{s}$. They separately act as the key and value of the text cross-attention layer in two branches.
The seed point features $V$ act as the key and value of the visual cross-attention layer in both two branches. Note that we design different adaptive position encoding methods for each branch, which will be described in detail in Sec. Adaptive Position Encoding. Finally, visual-guided and text-guided features in each branch are fused together first, and then features from the two branches are fused. The query proposal features are updated as $V^{h}_{q}\in\mathbb{R}^{K\times D}$, where $h$ represents the $h-th$ layer of the decoder.
Figure 3: Different methods to calculate the relevant position between the sample of seed points A and the predicted box: (a) Vertex Relative Positioning (Shen et al. 2024); (b) Center Relative Positioning; (c)(d) Box-surface Relative Positioning for points outside/inside the box.
@@
Following (Wu et al. 2023), the query proposal features $V^{h}_{q}$ in each layer are projected into position features $V^{h}_{l}\in\mathbb{R}^{K\times D_{l}}$ and object semantic features $V^{h}_{o}\in\mathbb{R}^{K\times D_{o}}$ for losses calculation, where $D_{l}$ and $D_{o}$ represents the dimensions of two types of features. A box prediction head (Liu et al. 2021) is employed to generate bounding boxes $B_{out}\in\mathbb{R}^{K_{out}\times 6}$.
Adaptive Position Encoding
Position Encoding for Attention Map.
Position encoding (PE) can effectively refine the attention map, which is crucial for understanding the contextual information of points in 3D scenes. For the 3D visual grounding task, we update the calculation method of the attention map for the decoder as follows:
@@
We encode the relative positions between seed points and coarsely predicted boxes as follows:
${\mathcal{E}}=\operatorname{MLP}(\operatorname{F}(\Delta\mathcal{E}),$ (2)
where $\operatorname{F}(\cdot)$ is a non-linear function and $\Delta\mathcal{E}\in\mathbb{R}^{K\times N\times 3}$ is the calculated relative position. $\operatorname{MLP}$ represents an MLP-based transformation that projects the features to a higher dimension space. For a detailed explanation of this part, please refer to (Shen et al. 2024).
@@
To represent the relative position between $N$ seed points and $K$ boxes, we find the point on the box surface closest to each seed point and use the offsets in three dimensions as $\Delta\mathcal{E}=\{(e^{ij}_{x},e^{ij}_{y},e^{ij}_{z})|\quad i=0,1,...,K;j=0,1,...,N\}$, which is shown in Fig. 3 (c)(d). To be specific, for one of seed points $A=(x_{0},y_{0},z_{0})$, our goal is to find a point $M=(x,y,z)$ on the box surface to minimize the distance between $M$ and $A$. The offset can be set as
$\begin{split}(e_{x},e_{y},e_{z})=(x_{0}-x,y_{0}-y,z_{0}-z).\end{split}$ (3)
We set no absolute values here in order to distinguish between points inside and outside the box with the same offset.
For a point $A$ outside the box as Fig. 3 (c), with a box $b=(x_{c},y_{c},z_{c},l,w,h)$ the three-dimensional offset can be represented as
$\begin{split}(e_{x},e_{y},e_{z})=(\operatorname{max}(|x_{0}-x_{c}|-\frac{l}{2},0),\\ \operatorname{max}(|y_{0}-y_{c}|-\frac{w}{2},0),\operatorname{max}(|z_{0}-z_{c}|-\frac{h}{2},0)).\end{split}$ (4)
When the seed point $A^{\prime}=(x^{\prime}_{0},y^{\prime}_{0},z^{\prime}_{0})$ is inside the box as Fig. 3 (d), the minimum distance of $M$ and $A^{\prime}$ can be represented as:
$MA^{\prime}_{min}=\operatorname{min}(\frac{l}{2}-|x^{\prime}_{0}-x_{c}|,\frac{w}{2}-|y^{\prime}_{0}-y_{c}|,\frac{h}{2}-|z^{\prime}_{0}-z_{c}|).$ (5)
If $MA^{\prime}$ reaches its minimum value in the x-direction, then $e_{x}$ is set to $-MA^{\prime}$, and the other two offsets are set to 0 because $M$ and $A^{\prime}$ have the same coordinate values in the y-direction and z-direction; the same applies to $e_{y}$ and $e_{z}$. The offset takes the negative of the distance to differentiate from the seed point outside the box. For detailed mathematical reasoning, please refer to the supplementary material.
@@
Compared to the method shown in Fig. 3 (b) that directly uses the offset between the point and the box center as the relative position, our Box-surface Relative Positioning method incorporates both relative distance information and the shape information of the box. In contrast to the Vertex Relative Positioning method proposed by (Shen et al. 2024), as shown in Fig. 3 (a), which uses eight MLP-based transformations in each layer, our method encodes only once per layer of cross-attention, making it more time-efficient and memory-saving. For detailed ablation experiments, please refer to Sec. Experiments.
Figure 4: The illustration of the Adaptive Position Encoding method. Step 3 is the text-guided confidence that is only used in the surrounding branch.
Unique ($\sim$19%) Multiple ($\sim$81%) Overall
Method Venue @0.25 @0.5 @0.25 @0.5 @0.25 @0.5
ScanRefer (Chen, Chang, and Nießner 2020) ECCV2020 67.6 46.2 32.1 21.3 39.0 26.1
TGNN (Huang et al. 2021) AAAI2021 68.6 56.8 29.8 23.2 37.4 29.7
InstanceRefer (Yuan et al. 2021) ICCV2021 77.5 66.8 31.3 24.8 40.2 32.9
SAT (Yang et al. 2021) ICCV2021 73.2 50.8 37.6 25.2 44.5 30.1
FFL-3DOG (Feng et al. 2021) ICCV2021 78.8 67.9 35.2 25.7 41.3 34.0
3DVG-Transformer (Zhao et al. 2021) ICCV2021 77.2 58.5 38.4 28.7 45.9 34.5
3D-SPS (Luo et al. 2022) CVPR2022 84.1 66.7 40.3 29.8 48.8 37.0
3DJCG (Cai et al. 2022) CVPR2022 78.8 61.3 40.1 30.1 47.6 36.1
BUTD-DETR (Jain et al. 2022) ECCV2022 82.9 65.0 44.7 34.0 50.4 38.6
D3Net (Chen et al. 2022a) ECCV2022 - 70.4 - 30.5 - 37.9
EDA (Wu et al. 2023) CVPR2023 85.8 68.6 49.1 37.6 54.6 42.3
ViewRefer (Guo et al. 2023) ICCV2023 - - 33.1 26.5 41.3 33.7
3DRP-Net (Wang et al. 2023) EMNLP2023 83.1 67.7 42.1 32.0 50.1 38.9
CORE-3DVG (Yang et al. 2023) NIPS2023 85.0 67.1 51.8 39.8 56.8 43.8
3DVLP (Zhang et al. 2023) AAAI2024 85.2 70.0 43.7 33.4 51.7 40.5
VPP-Net (Shi, Wu, and Lee 2024) CVPR2024 86.1 67.1 50.3 39.0 55.7 43.3
PD-APE (ours-PointNet++) 86.7 73.2 50.2 40.6 56.6 45.5
PD-APE (ours-Sparse3D) 87.7 75.1 52.6 42.2 57.8 47.1